# Distributed Constrained Online Nonconvex Optimization with Compressed Communication

### v1

@@ Abstract
This paper considers distributed online nonconvex optimization with time-varying inequality constraints over a network of agents. For a time-varying graph, we propose a distributed online primal–dual algorithm with compressed communication to efficiently utilize communication resources. We show that the proposed algorithm establishes an $\mathcal{O}({{T^{\max\{{1-{\theta_{1}},{\theta_{1}}}\}}}})$ network regret bound and an $\mathcal{O}({T^{1-{\theta_{1}}/2}})$ network cumulative constraint violation bound, where $T$ is the number of iterations and ${\theta_{1}}\in({0,1})$ is a user-defined trade-off parameter. When Slater’s condition holds (i.e, there is a point that strictly satisfies the inequality constraints at all iterations), the network cumulative constraint violation bound is reduced to $\mathcal{O}({T^{1-{\theta_{1}}}})$. These bounds are comparable to the state-of-the-art results established by existing distributed online algorithms with perfect communication for distributed online convex optimization with (time-varying) inequality constraints. Finally, a simulation example is presented to validate the theoretical results.
@@ I Introduction
Distributed online convex optimization offers a promising framework for modeling a variety of problems in dynamic, uncertain, and adversarial environments, with wide-ranging applications such as real-time routing in data networks and online advertisement placement in web search [1]. This framework can be understood as a structured repeated game with $T$ iterations between a network of agents and an adversary. Specifically, at each iteration $t$, each agent $i$ selects a decision ${x_{i,t}}\in\mathbb{X}$, where $\mathbb{X}\subseteq{\mathbb{R}^{p}}$ is a known convex set and $p$ is a positive integer. Upon selection, the local loss function ${f_{i,t}}$ is privately revealed to agent $i$ by the adversary. The goal of agents is to collaboratively minimize the network-wide accumulated loss, and the corresponding performance metric is network regret
$\frac{1}{n}\sum\limits_{i=1}^{n}{\Big{(}{\sum\limits_{t=1}^{T}{{f_{t}}({{x_{i,t}}})}-\mathop{\min}\limits_{x\in\mathbb{X}}\sum\limits_{t=1}^{T}{{f_{t}}({{x}})}}\Big{)}},$
where ${f_{t}}(x):=\frac{1}{n}\sum\nolimits_{j=1}^{n}{{f_{j,t}}(x)}$ is the global loss function of the network at iteration $t$.
@@
Various projection-based distributed online algorithms with sublinear regret have been proposed to solve the distributed online convex optimization problem, see, e.g., [2, 3, 4, 5, 6, 7, 8, 9, 10], and recent survey paper [11]. For example, for the fix communication topology, the authors of [2] propose a projection-based distributed online subgradient descent algorithm, and establish an $\mathcal{O}({\sqrt{T}})$ regret bound for general convex local loss functions. For strongly convex local loss functions, the authors of [3, 4] establish an $\mathcal{O}\big{(}{\log(T)}\big{)}$ regret bound. For time-varying communication topology, the authors of [5] propose a projection-based distributed online weighted dual averaging algorithm, and establish an $\mathcal{O}({\sqrt{T}})$ regret bound for general convex local loss functions. By utilizing proportional-integral distributed feedback on the disagreement among neighboring agents, the authors of [6] propose a projection-based distributed online proportional-integral subgradient descent algorithm, and establish an $\mathcal{O}\big{(}{\log(T)}\big{)}$ regret bound for strongly convex local loss functions.
@@
The aforementioned distributed online algorithms rely on agents exchanging their local data with perfect communication. Consequently, these algorithms encounter significant limitations arising from communication bottlenecks, particularly in scenarios involving limited bandwidth and power resources. The limitations becomes more pronounced as the scale of the multi-agent network and the dimensionality of the exchanged data increase. To overcome the limitations, distributed online algorithms with compressed communication are studied for the fix communication topology in the literature, see [12, 13, 14], and recent survey paper [15]. For example, the authors of [12] propose a decentralized online gradient descent algorithm with compressed communication by introducing an auxiliary variable to estimate the neighbors’ decisions at each iteration. They establish an $\mathcal{O}(\sqrt{T})$ network regret bound for general convex local loss functions. Unlike the compression strategy employed in [12], the authors of [13, 14] introduce two auxiliary variables: the first serves the same purpose as the auxiliary variable in [12], while the second ensures that the first variable does not need to be exchanged. The compression strategy is effective when the communication topology is fixed. However, it becomes ineffective for a time-varying communication topology.
@@
Notations: All inequalities and equalities throughout this paper are understood componentwise. ${\mathbb{N}_{+}}$, $\mathbb{R}$, ${\mathbb{R}^{p}}$ and $\mathbb{R}_{+}^{p}$ denote the sets of all positive integers, real numbers, $p$-dimensional and nonnegative vectors, respectively. Given $m$ and $n\in{\mathbb{N}_{+}}$, $[m]$ denotes the set $\{{1,\cdot\cdot\cdot,m}\}$, and $[m,n]$ denotes the set $\{{m,\cdot\cdot\cdot,n}\}$ for $m<n$. Given vectors $x$ and $y$, ${x^{T}}$ denotes the transpose of the vector $x$, and $\langle{x,y}\rangle$ denotes the standard inner. ${\mathbf{0}_{p}}$ and ${\mathbf{1}_{p}}$ denote the $p$-dimensional column vector whose components are all 00 and $1$, respectively. $\mathrm{col}({q_{1}},\cdot\cdot\cdot,{q_{n}})$ denotes the concatenated column vector of ${q_{i}}\in{\mathbb{R}^{{m_{i}}}}$ for $i\in[n]$. ${\mathbb{B}^{p}}$ and ${\mathbb{S}^{p}}$ denote the unit ball and sphere centered around the origin in ${\mathbb{R}^{p}}$ under Euclidean norm, respectively. For a set $\mathbb{K}\in{\mathbb{R}^{p}}$ and a vector $x\in{\mathbb{R}^{p}}$, ${\mathcal{P}_{\mathbb{K}}}(x)$ denotes the projection of the vector $x$ onto the set $\mathbb{K}$, i.e., ${\mathcal{P}_{\mathbb{K}}}(x)=\arg{\min_{y\in{\mathbb{K}}}}{\|{x-y}\|^{2}}$, and $[x]_{+}$ denotes ${\mathcal{P}_{\mathbb{R}_{+}^{p}}}(x)$. For a function $f$ and a vector $x$, $\nabla f(x)$ denotes the subgradient of $f$ at $x$.
@@ II Problem Formulation
For all $i\in[n]$, $t\in{\mathbb{N}_{+}}$, there exists a constant ${L}$ such that
@@
Remark 1.
Note that we do not require the assumptions that the local loss function ${{f_{i,t}}}$ and local constraint function ${{g_{i,t}}}$ are uniformly bounded, whereas [20] imposes them, and [24] imposes the first one. In addition, Assumption 2 is commonly used for distributed online convex optimization with long-term constraints [16, 18, 20, 24]. Since the local constraint function ${{g_{i,t}}}$ is convex, Assumption 2 implies its Lipschitz continuity on $\mathbb{X}$. In contrast, the nonconvex nature of the local constraint function ${{f_{i,t}}}$ precludes the existence of such a Lipschitz continuity property. This fundamental distinction causes analytical challenges compared to distributed online convex optimization, particularly in establishing network regret bounds. Furthermore, Assumption 3 plays a crucial role in addressing the analytical challenges, which is also used in distributed online nonconvex optimization [25].
Different from the studies on distributed online convex optimization with compressed communication [27, 12, 13, 14] that consider a fix and undirected graph, we consider a time-varying directed graph. The following assumption on the graph is made, which is also used in [16, 20, 24].
@@
In this paper, we consider the scenario where the communication between agents is compressed by using a class of compressors with globally bounded absolute compression error, as given in the following.
@@
The compressor $\mathcal{C}:{\mathbb{Y}}\to{\mathbb{Y}}$ with $\mathbb{Y}\in{\mathbb{R}^{p}}$ satisfies
${\mathbf{E}_{\mathcal{C}}}[{\|{\mathcal{C}(x)-x}\|_{d}^{2}}]\leq C,\forall x\in{\mathbb{Y}},$ (7)
@@
Assumption 5 covers the majority of popular compressors in machine learning and signal processing applications such as the deterministic quantization used in [28, 29] and the unbiased stochastic quantization used in [30, 31]. The same class of compressors satisfying Assumption 5 is also used in [32, 14, 33].
Next, Slater’s condition is present in the following.
@@
Slater’s condition is a sufficient condition for strong duality to hold in convex optimization problems [21]. However, for nonconvex problems, Slater’s condition alone does not generally guarantee strong duality because the standard Lagrange dual often has a nonzero duality gap in nonconvex problems [34]. For distributed online convex optimization, [24] establishes reduced network cumulative constraint violation bounds under Slater’s condition. However, to the best of our knowledge, there are no studies to show that similar reductions for cumulative constraint violation bounds can be achieved in distributed online nonconvex optimization, even for the general constraint violation bounds.
@@ III Distributed Online Primal–Dual Algorithm with
In this section, we propose a distributed online primal–dual algorithm with compressed communication by using the class of compressors satisfying Assumption 5, and analyze the performance of the algorithm without and with Slater’s condition, respectively.
@@ III-A Algorithm Description
where ${{\gamma_{t}}}$ is the regularization parameter; and ${{\alpha_{t}}}$ is the stepsize.
To implement the algorithm (9), at each iteration each agent $i$ need to receive the exact vector-valued variable $z_{j,t}$ from its neighbors. That results in a substantial amount of data exchange over all iterations, especially when the scale of the multi-agent network and the dimension $p$ are large. To improve communication efficiency, we use the compressor satisfying Assumption 5 to compress $z_{j,t}$. Note that directly using the compressed variable $\mathcal{C}({{z_{j,t}}})$ to replace $z_{j,t}$ in (9a) does not work due to the compression error. To reduce the compression error, an auxiliary variable $\hat{z}_{j,t}\in{\mathbb{R}^{p}}$ and a scaling parameter ${s_{t}}$ can be introduced. Instead of directly compressing $z_{j,t}$, one can compress the scaled difference ${({{z_{j,t}}-{\hat{z}_{j,t-1}}})/{s_{t}}}$, then multiply the compressed result by ${s_{t}}$ and add it back to ${{\hat{z}}_{j,t-1}}$ to replace $z_{j,t}$ in (9a). Thus, (9a) is modified as follows:
${x_{i,t}}=\sum\limits_{j=1}^{n}{{{[{W_{t}}]}_{ij}}{\hat{z}_{j,t}}},$ (10)
where
${\hat{z}_{j,t}}$ $={\hat{z}_{j,t-1}}+{s_{t}}\mathcal{C}\big{(}{({{z_{j,t}}-{\hat{z}_{j,t-1}}})/{s_{t}}}\big{)}.$ (11)
In this way, the compression error is reduced. However, as stated in [32], at each iteration each agent $i$ needs to receive the exact vector-valued variable $\hat{z}_{j,t-1}$ from its neighbors due to the term $\sum\nolimits_{j=1}^{n}{{{[{W_{t}}]}_{ij}}{{\hat{z}}_{j,t-1}}}$ in (10). Therefore, the improved algorithm does not enjoy the benefits of compression.
To deal with the dilemma, [32] introduces another auxiliary variable ${y_{j,t}}\in{\mathbb{R}^{p}}$ to remove the term $\sum\nolimits_{j=1}^{n}{{{[{W_{t}}]}_{ij}}{{\hat{z}}_{j,t-1}}}$. Then, (9a) becomes the following form:
${x_{i,t}}={{\hat{z}}_{j,t}}-{y_{j,t}},$ (12)
where
${{\hat{z}}_{j,t}}$ $={{\hat{z}}_{j,t-1}}+{s_{t}}{\cal C}\big{(}({z_{j,t}}-{{\hat{z}}_{j,t-1}})/{s_{t}}\big{)},$ (13a)
${y_{j,t}}$ $={y_{j,t-1}}+{s_{t}}{\cal C}\big{(}({z_{j,t}}-{{\hat{z}}_{j,t-1}})/{s_{t}}\big{)}-{s_{t}}\sum\limits_{j=1}^{n}{{{[{W_{t}}]}_{ij}}{\cal C}\big{(}({z_{j,t}}-{{\hat{z}}_{j,t-1}})/{s_{t}}\big{)}}.$ (13b)
Note that agent $i$ only requires the compressed data as presented in (13). The similar strategy is also used in [13, 14]. When the graph $\mathcal{G}_{t}$ is fixed, i.e., ${[{W_{t-1}}]_{ij}}={[{W_{t}}]_{ij}}$ for all $t\in[T]$, it is straightforward to check that ${y_{j,t}}={{\hat{z}}_{j,t}}-\sum\nolimits_{j=1}^{n}{{{[{W_{t}}]}_{ij}}{{\hat{z}}_{j,t}}}$ for $j\in[n]$ by mathematical induction. However, the critical result does not hold when the graph $\mathcal{G}_{t}$ is time-varying.
To handle this challenge, instead of introducing another auxiliary variable, we define $\hat{z}_{j,t}$ in (10) as a vector-valued variable stored in agent $i$ rather than its neighbors. The variable $\hat{z}_{j,t}$ is indexed by $j$ to maintain one-to-one correspondence with all agents. Consequently, its value no longer needs to be exchanged, and only the compressed data $\mathcal{C}({({{z_{j,t}}-{{\hat{z}}_{j,t-1}}})/{s_{t}}})$ in (11) is required by agent $i$. In this way, the distributed online primal–dual algorithm with compressed communication is proposed, which is presented in pseudo-code as Algorithm 1.
@@
${\hat{z}_{j,t}}$ $={\hat{z}_{j,t-1}}+{s_{t}}\mathcal{C}\big{(}{({{z_{j,t}}-{\hat{z}_{j,t-1}}})/{s_{t}}}\big{)},j\in\big{\{}\mathcal{N}_{i}^{\text{in}}({{\mathcal{G}_{t}}})\cup\{i\}\big{\}}.$ (14)
@@
It is worth noting that agent $j$ does not need to receive ${{\hat{z}}_{j,t-1}}$ in (14) from any other agents to compute $\mathcal{C}({({{z_{j,t}}-{{\hat{z}}_{j,t-1}}})/{s_{t}}})$, as it naturally possesses this information. In Algorithm 1, we initialize ${\hat{z}_{j,0}}={\mathbf{0}_{p}}$ for $j\in[n]$, and thus ${\hat{z}_{j,1}}={s_{1}}{\cal C}({z_{j,1}}/{s_{1}})$, which is inherently known by agent $j$ since $z_{j,1}$ is precisely stored in agent $j$. More generally, at each iteration $t$, agent $j$ has access to both $z_{j,t}$ and $\hat{z}_{j,t-1}$, enabling it to independently determine $\hat{z}_{j,t}$ without requiring information from any other agents.
@@ III-B Performance Analysis
Due to the space limitations, we omit the proof, which can be found in the arXiv version
@@
We show in Theorem 1 that Algorithm 1 establishes sublinear network regret and cumulative constraint violation bounds as in (20)–(21). These bounds characterize the impact of compressed communication on the network regret and cumulative constraint violation bounds, which is captured by ${\theta_{2}}$. When ${\theta_{2}}\geq 1$, compressed communication does not affect either of the bounds, and they are the same as the state-of-the-art results established by the distributed online algorithms without compressed communication in [20, 24]. When ${\theta_{1}}<{\theta_{2}}<1$, compressed communication may enable the network regret bound to become larger due to $01-{\theta_{2}}>$ 0, but does not affect the network cumulative constraint violation bound due to ${\theta_{2}}>{\theta_{1}}$. In addition, when Slater’s condition holds, the network cumulative constraint violation bound is further reduced as in (22). The bound is unaffected by compressed communication and remains the same as the results established by the distributed online algorithm with perfect communication in [24].
@@
Due to the space limitations, we omit the proof, which can be found in the arXiv version
@@ IV Simulation Example
$\;{\rm{s}}{\rm{.t}}{\rm{.}}\;\;\;\;x\in\mathbb{X},{\rm{}}{B_{i,t}}x-{b_{i,t}}\leq{\mathbf{0}_{{m_{i}}}},\forall i\in[n],\forall t\in[T].$ (27b)
In this problem, the sensors aim to cooperatively track a moving target. The position of sensor $i$ is denoted by $S_{i}\in{\mathbb{R}^{p}}$. Sensor $i$ measures the distance between the positions of the target and itself by ${D_{i,t}}={\|{{S_{i}}-{X_{0,t}}}\|^{2}}+{\tau_{i,t}}$ where ${X_{0,t}}\in{\mathbb{R}^{p}}$ and ${\tau_{i,t}}\in\mathbb{R}$ denote the position of the target and the measurement noise at iteration $t$, respectively. Moreover, each agent only has access to its distance measurement. In addition, the position of the target is required to remain within a designated safe region characterized by the set $\mathbb{X}$, and avoid prolonged deviations from the mission region characterized by time-varying linear inequality constraint ${B_{i,t}}x-{b_{i,t}}\leq{\mathbf{0}_{{m_{i}}}}$ with ${B_{i,t}}\in{\mathbb{R}^{{m_{i}}\times p}}$ and ${b_{i,t}}\in{\mathbb{R}^{{m_{i}}}}$. However, occasional deviations are permitted to accommodate obstacle avoidance and exploration requirements. Inspired by [35, 36, 37], the problem as in (27) is formulated to achieve the least squares estimator for the position of the target. The communication topology is modeled by a time-varying undirected graph. Specifically, at each iteration $t$, the graph is first randomly generated where the probability of any two sensors being connected is $\rho$. Then, to make sure that Assumption 4 is satisfied, we add edges $({i,i+1})$ for $i\in[24]$ when $t\in\{{4c+1}\}$, edges $({i,i+1})$ for $i\in[25,49]$ when $t\in\{{4c+2}\}$, edges $({i,i+1})$ for $i\in[50,74]$ when $t\in\{{4c+3}\}$, edges $({i,i+1})$ for $i\in[75,99]$ when $t\in\{{4c+4}\}$ for $c=\{0,1,\cdot\cdot\cdot\}$. Moreover, let ${[{{W_{t}}}]_{ij}}=\frac{1}{n}$ if $({j,i})\in{\mathcal{E}_{t}}$ and ${[{{W_{t}}}]_{ii}}=1-\sum\nolimits_{j=1}^{n}{{{[{{W_{t}}}]}_{ij}}}$.
In this paper, we show in Theorem 1 that, both without and with Slater’s condition, Algorithm 1 establishes the same network regret and cumulative constraint violation bounds as the state-of-the-art results on distributed online convex optimization with long-term constraints, established by the distributed online algorithms with perfect communication in [24]. To verify the theoretical results, we compare Algorithm 1 with the algorithm in [24]. We set $\rho=0.1$, $\mathbb{X}={[{-5,5}]^{p}}$, $p=2$, ${m_{i}}=2$, and randomly choose each component of ${{S_{i}}}$ from the uniform distribution in the interval $[{-10,10}]$. We assume that the position of the target evolves by
@@
where ${{Q_{t}}}$ is randomly generated from Bernoulli distribution with a success probability of $0.5$, and ${X_{0,0}}={[{0.8,0.95}]^{T}}$. Moreover, ${\tau_{i,t}}$ is randomly generated from the uniform distribution from in the interval $[{0,0.001}]$. Furthermore, each component of ${B_{i,t}}$ is randomly generated from the uniform distribution in the interval $[{0,2}]$, and each component of ${b_{i,t}}$ is randomly generated from the uniform distribution in the interval $[{b,b+1}]$ with $b>0$. Note that $b>0$ guarantees Slater’s condition holds. Here we choose $b=0.01$. In addition, we select the following compressor for Algorithm 1.
Standard uniform quantizer:
@@
where $\Delta$ is a positive integer. This compressor satisfies Assumption 5 with $d=\infty$ and $C={\Delta^{2}}/4$, which is also used in [32, 14, 29, 33]. Transmitting $\mathcal{C}(x)$ requires $pq$ bits if each integer is encoded using $q$ bits. Here we set $\Delta=1$ and $q=8$. The inputs of all these algorithms are listed in TABLE I.
TABLE I: Input of algorithms.
Algorithms Inputs
Algorithm 1 in this paper ${\alpha_{t}}=0.5/t,{\gamma_{t}}=0.15/{\alpha_{t}},{s_{t}}=1/t$
The algorithm in [24] ${\alpha_{t}}=0.5/t,{\gamma_{t}}=0.15/{\alpha_{t}},\psi(x)={\|x\|^{2}}/2$

### v2

@@ Abstract
This paper considers distributed online nonconvex optimization with time-varying inequality constraints over a network of agents. For a time-varying graph, we propose a distributed online primal–dual algorithm with compressed communication to efficiently utilize communication resources. We show that the proposed algorithm establishes an $\mathcal{O}({{T^{\max\{{1-{\theta_{1}},{\theta_{1}}}\}}}})$ network regret bound and an $\mathcal{O}({T^{1-{\theta_{1}}/2}})$ network cumulative constraint violation bound, where $T$ is the number of iterations and ${\theta_{1}}\in({0,1})$ is a user-defined trade-off parameter. When Slater’s condition holds, the network cumulative constraint violation bound is reduced to $\mathcal{O}({T^{1-{\theta_{1}}}})$. These bounds are comparable to the state-of-the-art results established by existing distributed online algorithms with perfect communication for distributed online convex optimization with (time-varying) inequality constraints. Finally, a simulation example is presented to validate the theoretical results.
@@ I Introduction
Distributed online convex optimization offers a promising framework for modeling a variety of problems in dynamic, uncertain, and adversarial environments, with wide-ranging applications such as real-time routing in data networks and online advertisement placement in web search [1]. Various projection-based distributed online algorithms with sublinear regret have been proposed to solve the distributed online convex optimization problem, see, e.g., [2, 3, 4, 5, 6, 7, 8, 9, 10], and recent survey paper [11]. For example, for the fix communication topology, the authors of [2] propose a projection-based distributed online subgradient descent algorithm, and establish an $\mathcal{O}({\sqrt{T}})$ regret bound for general convex local loss functions. For strongly convex local loss functions, the authors of [3, 4] establish an $\mathcal{O}\big{(}{\log(T)}\big{)}$ regret bound. For time-varying communication topology, the authors of [5] propose a projection-based distributed online weighted dual averaging algorithm, and establish an $\mathcal{O}({\sqrt{T}})$ regret bound for general convex local loss functions. By utilizing proportional-integral distributed feedback on the disagreement among neighboring agents, the authors of [6] propose a projection-based distributed online proportional-integral subgradient descent algorithm, and establish an $\mathcal{O}\big{(}{\log(T)}\big{)}$ regret bound for strongly convex local loss functions.
@@
The aforementioned distributed online algorithms rely on agents exchanging their local data with perfect communication. Consequently, these algorithms encounter significant limitations arising from communication bottlenecks. To overcome the limitations, distributed online algorithms with compressed communication are studied for the fix communication topology in the literature, see [12, 13, 14], and recent survey paper [15]. For example, the authors of [12] propose a decentralized online gradient descent algorithm with compressed communication by introducing an auxiliary variable to estimate the neighbors’ decisions at each iteration. They establish an $\mathcal{O}(\sqrt{T})$ network regret bound for general convex local loss functions. Unlike the compression strategy employed in [12], the authors of [13, 14] introduce two auxiliary variables: the first serves the same purpose as the auxiliary variable in [12], while the second ensures that the first variable does not need to be exchanged. The compression strategy is effective when the communication topology is fixed. However, it becomes ineffective for a time-varying communication topology.
@@
Notations: All inequalities and equalities throughout this paper are understood componentwise. $\mathbb{N}_{+}$}, $\mathbb{R}$, $\mathbb{R}^{p}$} and $\mathbb{R}_{+}^{p}$ denote the sets of all positive integers, real numbers, $p$-dimensional and nonnegative vectors, respectively. Let $\mathbb{X}$ denote a subset of $\mathbb{R}^{p}$}. Given $m$ and $n\in{\mathbb{N}_{+}}$, $[m]$ denotes the set $\{{1,\cdot\cdot\cdot,m}\}$, and $[m,n]$ denotes the set $\{{m,\cdot\cdot\cdot,n}\}$ for $m<n$. Given vectors $x$ and $y$, $x^{T}$} denotes the transpose of the vector $x$, and $\langle{x,y}\rangle$ denotes the standard inner. $\mathbf{0}_{p}$} and $\mathbf{1}_{p}$} denote the $p$-dimensional column vector whose components are all 0 and $1$, respectively. $\mathrm{col}({q_{1}},\cdot\cdot\cdot,{q_{n}})$ denotes the concatenated column vector of ${q_{i}}\in{\mathbb{R}^{{m_{i}}}}$ for $i\in[n]$. $\mathbb{B}^{p}$} and $\mathbb{S}^{p}$} denote the unit ball and sphere centered around the origin in $\mathbb{R}^{p}$} under Euclidean norm, respectively. For a set $\mathbb{K}\in{\mathbb{R}^{p}}$ and a vector $x\in{\mathbb{R}^{p}}$, ${\mathcal{P}_{\mathbb{K}}}(x)$ denotes the projection of the vector $x$ onto the set $\mathbb{K}$, i.e., ${\mathcal{P}_{\mathbb{K}}}(x)=\arg{\min_{y\in{\mathbb{K}}}}{\|{x-y}\|^{2}}$, and $x]_{+}$ denotes ${\mathcal{P}_{\mathbb{R}_{+}^{p}}}(x)$. For a function $f$ and a vector $x$, $\nabla f(x)$ denotes the subgradient of $f$ at $x$.
@@ II Problem Formulation
For all $i\in[n]$, $t\in{\mathbb{N}_{+}}$, there exists a constant L{L} such that
@@
The compressor $\mathcal{C}:{\mathbb{R}^{p}}\to{\mathbb{R}^{p}}$ satisfies
${\mathbf{E}_{\mathcal{C}}}[{\|{\mathcal{C}(x)-x}\|_{d}^{2}}]\leq C,\forall x\in{\mathbb{R}^{p}},$ (7)
@@ III-A Algorithm Description
where $\gamma_{t}$}} is the regularization parameter; and $\alpha_{t}$}} is the stepsize. To improve communication efficiency, the distributed online primal–dual algorithm with compressed communication is proposed by using the class of compressors satisfying Assumption 5, which is presented in pseudo-code as Algorithm 1.
@@
${{\hat{z}}_{j,t}}={\mathcal{P}_{\mathbb{X}}}\big{(}{{{\hat{z}}_{j,t-1}}+{s_{t}}{\cal C}(({z_{j,t}}-{{\hat{z}}_{j,t-1}})/{s_{t}})}\big{)},j\in\big{\{}\mathcal{N}_{i}^{\text{in}}({{\mathcal{G}_{t}}})\cup\{i\}\big{\}}.$ (10)
@@ III-B Performance Analysis
We show in Theorem 1 that Algorithm 1 establishes sublinear network regret and cumulative constraint violation bounds as in (16)–(17). These bounds characterize the impact of compressed communication on the network regret and cumulative constraint violation bounds, which is captured by $\theta_{2}$}. When ${\theta_{2}}\geq 1$, they are the same as the state-of-the-art results established by the distributed online algorithms without compressed communication in [20, 24]. In addition, when Slater’s condition holds, the network cumulative constraint violation bound is further reduced as in (18). The bound remains the same as the results established by the distributed online algorithm with perfect communication in [24].
@@ IV Simulation Example
$\;{\rm{s}}{\rm{.t}}{\rm{.}}\;\;\;\;x\in\mathbb{X},{\rm{}}{B_{i,t}}x-{b_{i,t}}\leq{\mathbf{0}_{{m_{i}}}},\forall i\in[n],\forall t\in[T],$ (23b)
where $S_{i}\in{\mathbb{R}^{p}}$ denotes the position of sensor $i$, ${D_{i,t}}={\|{{S_{i}}-{X_{0,t}}}\|^{2}}+{\tau_{i,t}}$ denotes the distance between the positions of the target and sensor $i$, ${X_{0,t}}\in{\mathbb{R}^{p}}$ and $\tau_{i,t}$$\in\mathbb{R}$ denote the position of the target and the measurement noise at iteration $t$, respectively, and ${B_{i,t}}\in{\mathbb{R}^{{m_{i}}\times p}}$ and ${b_{i,t}}\in{\mathbb{R}^{{m_{i}}}}$ denote the coefficient matrix and coefficient vector of the local linear constraints, respectively. The communication topology is modeled by a time-varying undirected graph. Specifically, at each iteration $t$, the graph is first randomly generated where the probability of any two sensors being connected is $\rho$. Then, to make sure that Assumption 4 is satisfied, we add edges (i,i+1)({i,i+1}) for $i\in[24]$ when $t\in\{{4c+1}\}$, edges (i,i+1)({i,i+1}) for $i\in[25,49]$ when $t\in\{{4c+2}\}$, edges (i,i+1)({i,i+1}) for $i\in[50,74]$ when $t\in\{{4c+3}\}$, edges (i,i+1)({i,i+1}) for $i\in[75,99]$ when $t\in\{{4c+4}\}$ for $c=\{0,1,\cdot\cdot\cdot\}$. Moreover, let ${[{{W_{t}}}]_{ij}}=\frac{1}{n}$ if $({j,i})\in{\mathcal{E}_{t}}$ and ${[{{W_{t}}}]_{ii}}=1-\sum\nolimits_{j=1}^{n}{{{[{{W_{t}}}]}_{ij}}}$.
In this paper, we show in Theorem 1 that, both without and with Slater’s condition, Algorithm 1 establishes the same network regret and cumulative constraint violation bounds as the state-of-the-art results on distributed online convex optimization with long-term constraints, established by the distributed online algorithms with perfect communication in [24]. To verify the theoretical results, we compare Algorithm 1 with the algorithm in [24]. We set $\rho=0.1$, $\mathbb{X}={[{-5,5}]^{p}}$, $p=2$, ${m_{i}}=2$, and randomly choose each component of $S_{i}$}} from the uniform distribution in the interval [10,10][{-10,10}]. We assume that the position of the target evolves by
@@
where $Q_{t}$}} is randomly generated from Bernoulli distribution with a success probability of $0.5$, and ${X_{0,0}}={[{0.8,0.95}]^{T}}$. Moreover, $\tau_{i,t}$} is randomly generated from the uniform distribution from in the interval [0,0.001][{0,0.001}]. Furthermore, each component of $B_{i,t}$} is randomly generated from the uniform distribution in the interval [0,2][{0,2}], and each component of $b_{i,t}$} is randomly generated from the uniform distribution in the interval [b,b+1][{b,b+1}] with $b>0$. Note that $b>0$ guarantees Slater’s condition holds. Here we choose $b=0.01$. In addition, we select the following compressor for Algorithm 1:
@@
where $\Delta$ is a positive integer. This compressor satisfies Assumption 5 with $d=\infty$ and $C={\Delta^{2}}/4$, which is also used in [27, 14, 28, 29]. Transmitting $\mathcal{C}(x)$ requires $pq$ bits if each integer is encoded using $q$ bits. Here we set $\Delta=1$ and $q=8$.